Determine the concavity , y-intercept , x-intercepts and co-ordinates of vertex of the parabola 𝑓(π‘₯) = 5π‘₯ 2 βˆ’ π‘₯ βˆ’ 3.Β 


1
Expert's answer
2021-11-22T05:42:39-0500

To find concavity we should find the second derivative

"f'(x)=10x-1"

"f''(x)=10>0\implies"f(x) is concave upward on the whole domain


To find y-inercepts we should put x = 0

"f(0)=5*0^2-0-3=-3"

point (0, -3) is y-intercept


To find x-intercepts we should put y = 0

"y=0\implies 5x^2-x-3=0"

"D=1+60=61\implies x_1={\frac{1+\sqrt{61}} {10}}, x_2={\frac{1-\sqrt{61}} {10}}"

points "({\frac{1+\sqrt{61}} {10}}, 0), ({\frac{1-\sqrt{61}} {10}}, 0)" is x-intercepts


Let point (a, b) be a vertex, then

"a={\frac {-b} {2a}}={\frac 1 {10}}"

to find b we should find f(0.1)

"f(0.1)=5*0.1^2-0.1-3=-3.05"

point (0.1, -3.05) is a vertex of the parabola